OpenStudy (anonymous):

piecewise functions... f(x)= x^2+2x if x is less than or equal to -1 x if -1is less than or equal to x less than or equal to1 -1 if x greater than 1

OpenStudy (anonymous):

what are we supposed to do with the piecewise function?

OpenStudy (anonymous):

evaluate the piecewise defined function at the indicated values.

OpenStudy (anonymous):

what are the indicated values? did they give any values to x?

OpenStudy (akshay_budhkar):

do you want to find whether its continuous and differentiable??

OpenStudy (anonymous):

neither.... i think its less simple than that... think of it as basic precal.

OpenStudy (anonymous):

that's all the book gives me....

OpenStudy (anonymous):

hmm...well usually it will say something like, "find the value of the piecewise function at x = -5" So then we have to figure out what interval -5 is in, and apply the correct piece of the function. Since -5 < -1, the correct piece is the x^2+2x and plugging in -5 gives: \[(-5)^2+2(-5) = 25-10 = 15\]

OpenStudy (anonymous):

If the question had asked what is the value at x = 0, we would use the function x since -1 < 0 < 1